OpenStudy (cutepuppies):

Which is the best estimate for the quotient of 404.31 ÷ 48.031? A. 8,000 B. 800 C. 80 D. 8

OpenStudy (cutepuppies):

@stefancipe

OpenStudy (retireed):

Estimate..... 400 / 50 = 40 / 5 = ________ ?

OpenStudy (anonymous):

just divide 404 with 48, and you'll see what is the best estimate :)